OpenStudy (calculator):

A buffer solution contains 0.1 M CH3COOH and 0.15M CH3COONa. 1.0cm3 of 0.1M NaOH was added to 1.0dm3 of the buffer solution.

OpenStudy (calculator):

What is the change in the pH of the solution? (Ka of CH3COOH = 1.75 x10-5 mol dm-3)

OpenStudy (calculator):

I should find the conc. of H+ right?

OpenStudy (calculator):

The previous question is linked to this question actually

OpenStudy (chmvijay):

i cant remeber that formula :(

OpenStudy (calculator):

Could it be \[CH_3COOH \rightleftharpoons CH_3COO^-+H^+\]

OpenStudy (calculator):

\[n(NaOH)=\frac{MV}{1000}=\frac{0.1}{1000}=1 \times 10^{-4}mol\]

OpenStudy (chmvijay):

i am sorry really i dont want to tell you the wrong one :) i really not getting it

OpenStudy (calculator):

Its alright I'll figure out on my own then, thanks

OpenStudy (chmvijay):

ur welcome

OpenStudy (chmvijay):

for ur last question this what happens when u add acid and base Consider a solution containing both acetic acid, CH3COOH, and acetate ions, CH3COO-. Any strong base that is added to the solution is neutralized by acetic acid: CH3COOH (aq) + OH-(aq) CH3COO-(aq) + H2O (aq) Any strong acid that is added to the solution is neutralized by acetate: CH3COO-(aq) + H+(aq) CH3COOH (aq) The amount of strong acid or base that a buffer can neutralize is called the buffer capacity. After the strong base or acid is neutralized, equilibrium will be reestablished.

OpenStudy (calculator):

Got it

OpenStudy (calculator):

Its done here :) http://answers.yahoo.com/question/index?qid=20130926060011AAkUcJH